Please show your steps in the calculations

6) Calculate the missing amounts: (Note: Year 1 End. R/E will be year 2 Beg. R/E Year Beg. R/E Balance Net income Dividends End. R/E Balance $2,100$950 300 2 3 2,600 2400800 d.
Answer #1
BEGINNING R/E BALANCE xxxxx
ADD: NET INCOME DURING THE YEAR xxxxx
LESS: DIVIDENDS DISTRIBUTE xxxxx
CLOSING R/E BALANCE xxxxxx
YEAR BEG. R/E BAL. (A) NET INCOME (B) DIVIDENDS (C) END. R/E BAL.(D)
A = LAST YEAR END. BALANCE B = D +C - A C = A + B - D D = A + B -C
1 $ 0 $ 2100 $ 950 $ 1150 (a)
2 $ 1150 (b) $ 1750 (c) $ 300 $ 2600
3. $ 2600 (d) $ 2400 $ 800 $ 4200 (e)